OpenStudy (anonymous):

A rocket sled for testing equipment under large accelerations starts at rest and acceler- ates according to the expression a= (1.5 m/s^3)t+ (2.9 m/s^2). How far does the sled move in the time interval t= 0 to t= 2.4 s?

OpenStudy (anonymous):

the integral of acceleration is velocity, then the integral of velocity is position, sound familiar?

OpenStudy (anonymous):

\[velocity = \int\limits_{ }^{ } acceleration = \int\limits_{ }^{ } 1.5t + 2.9 = 0.75t^{2} + 2.9t\]

OpenStudy (anonymous):

can you do the integral of 0.75t^2 + 2.9t to get the position equation, which then we can just plug in the time and get the answer?